# Regularity Priors for the Linear Atomic Cluster Expansion

### v1

@@ III.1 Silicon
In the next test an atom in the top layer of a Si slab, (110) surface, was moved in the $z$ direction until the atom sat just below the bottom surface of the slab due to periodic boundary conditions. The energy and z-component of the force on the target atom are shown in Figure 7, with, as before, increasing the regularity prior leading to better agreement with the Density-Functional Theory reference, particularly for the force. Notably, using larger $\sigma$ smooths out the unphysical oscillations present from 2 Å onwards but does not smooth out the relatively sharp feature around 0.5 Å which is present in the Density-Functional Theory reference. The same test was repeated for the (100) and (111) surfaces with the results shown in Figure S3.
@@
As a final test, the energy of various know Si phases, including diamond, bct5 and the simple-hexagonal (sh) polymorph, were computed as the volumes were varied isotropically. The resulting energy vs volume curves are provided in Figure 8 with pressure vs volume in Figure S3. Under compression, holes are present in the potentials with $\sigma=0$, 0.5 and 1.0, despite the dimer curves for $\sigma=0.5$ and $\sigma=1.0$ being repulsive at close-approach. Increasing $\sigma$ removes this issue and is a clear sign that the regularity prior is improving the behavior (and extrapolation) of the higher body-order terms. All ACE potentials correctly reproduce the curvature near the equilibrium volume whilst those with larger regularity priors capture more of the an-harmonic behavior at larger volumes. It is worth noting that the GAP and $\sigma=0$ potentials both exhibit additional minima in the energy-volume curves at large volumes. The large energy barriers to these states means they would not be encountered during typical MD simulations initialized near equilibrium conditions. However, they could cause issues in e.g. random structure search (see Figure 10) or nested sampling, which explore a much wider range of phase space.
@@
Next, ACE potentials were fit to the full Silicon-GAP-18 dataset. The aforementioned tests were repeated with the new models, see Figure S7, and similar improvements were seen when utilising a regularity prior, suggesting that such priors still improve performance even in a much more data-rich setting. Secondly, following [deringer2021origins], compression simulations were performed, wherein the low-density amorphous (LDA) structure of Si from [deringer2021origins] was gradually compressed during a molecular dynamics simulation by applying a pressure ramp from zero to 20 GPa. It is expected [deb2001pressure, mcmillan2005density, pandey2011pressure] that the material undergoes a sequence of phase transitions LDA $\rightarrow$ VHDA (very high density amorphous) $\rightarrow$ polycrystalline simple-hexagonal (pc-sh). This sequence is known to be reproduced by GAP models trained on the full Silicon-GAP-18 dataset [deringer2021origins], and can serve as a challenging test for interatomic potentials. [morrow2022indirect, morrow2023validate] Indeed, it was previously shown that the nucleation of pc-sh was not consistently produced by Moment Tensor Potentials (MTPs) trained directly on the Si-GAP-18 dataset [morrow2022indirect]. Instead, consistent nucleation could be achieved with MTP models if the training dataset was augmented with large, 1000 atom, configurations spanning the desired pressure range but labelled with a GAP model fit to the original data. This “student–teacher” model relationship was exploited to transfer the desired behaviour from the accurate, but slower GAP model to the much faster MTP, enabling larger simulations to be performed. In the present work, we explore whether using a regularity prior might allow us to bypass the need for the addition of synthetic data and instead to achieve the desired sequence of phase transitions with ACE models fitted directly to the Silicon-GAP-18 dataset.
@@
The results of such compression simulations, carried out as in ref. [morrow2022indirect], are shown in Figure 9, where the SOAP similarity to the pressure adjusted pc-sh phase is shown as a function of pressure for ACE models fit with 11 different levels of Gaussian regularity prior. With the default prior, $\sigma=0$, the dynamics become unstable during the transition from LDA $\rightarrow$ VHDA i.e. the model predicts spuriously large forces that accelerate atoms to velocities exceeding the integrator’s stability threshold, resulting in lost atoms and termination of the simulation. Prior to the crash, we observe the formation of chemically-unreasonable clusters of highly-coordinated Si, visible as blue patches in Figure 9B–C). These are false minima (“holes”) in the PES accessed under the influence of compression. Increasing $\sigma$ leads to a gradual increase in stability, with the $\sigma=1.0$ model successfully completing the full simulation. With moderate levels of regularity prior, $\sigma=1.0-2.0$, nucleation of the pc-sh phase is successfully observed, indicating that a well chosen regularity prior has helped mitigate the need for the student-teacher relationship. At larger values of $\sigma$ the dynamics becomes less stable, with no successful nucleation found to occur. This worsening performance correlates with the increasing errors shown in Figure S6, and implies that a well chosen $\sigma$ is one that approximately minimizes the force error on the test set.
@@
Figure 9: Following ref. [morrow2022indirect], the average SOAP similarity to the pressure adjusted sh-Si phase is shown for 100,000 atom simulations of amorphous Si during compression from 0-20 GPa. The sharp increase around 10 GPa indicates the transition from LDA $\rightarrow$ VHDA whilst the subsequent increase occurs on nucleation of the pc-sh phase. The difference in SOAP similarity compared to the GAP reference is shown below (note the break in the y-axis) whilst representative snapshots are highlighted above. A) Atoms are colored according to local coordination number to highlight an erroneously highly coordinated cluster in the top right corner which occurs due to a hole in the potential. B) and D) No nucleation of the sh-Si phase is seen with some regions adopting an unusual local structure with hints of hexagonal symmetry C) nucleation of the sh-crystal is seen with multiple differnet nucleation sites present. Similar structures are seen with $\sigma=1.41$ and 2.0Å, with simulations remaining stable as the growing grains impinge on each other, likely inducing large local stresses.
@@
Figure 10: The distribution of energy, relative to the diamond structure, and volume of random structures relaxed with various energy methods are shown. The axes are linear within the central square bounded by dashed grey lines and logarithmic outside. The bottom right panel shows four illustrative structures which are all predicted to be overly stable. a) consists of isolated clusters, b) is a 3D lattice where the closest distance between any pair of atoms is 3.7Å  and c) and d) are both 1D “nano-wires” where some atoms have more than 4 immediate neighbours.
@@
Following this, the Silicon ACE potentials were used to perform a random structure search. The search was restricted to 8 atom unit cells with 1000 putative structures generated without symmetry constraints and a further 2000 generated using 2-4 randomly chosen symmetry operations. Minimum distance constraints of 1.7Å  were enforced between neighbouring atoms and the volume of the initial cell was drawn from a uniform distribution spanning 12-30 $\text{\AA }^{3}$/atom. Structural relaxations were carried out with 0.01 GPa of external pressure applied in lieu of a dispersion-correction scheme and a force tolerance of 0.001 eV/Å. The distribution of relaxed structures is shown in Figure 10 for searches performed using Density-Functional Theory (CASTEP [clark2005first]), GAP and representative ACE models. The GAP search correctly identified the diamond structure as lowest in energy but contains a much larger proportion of low density, high energy structures than the Density-Functional Theory reference. In contrast, the ACE models all incorrectly find at least one structure to be lower in energy than diamond, with such structures all comprised of locally dense, highly coordinated, Si clusters or nano-wires. Among the ACE potentials, using a non-zero $\sigma$ significantly improves the marginal energy and volume distributions compared to the Density-Functional Theory reference and helps removes the low density, low energy structures, as might be expected from Figure 8 where the spurious oscillations in the volume distribution are eliminated. However, we stress that a regularity prior will not remove all spurious minima in the predicted PES and indeed this is the case, as is seen for the search performed with $\sigma=2.0$. Perhaps the clearest benefit of applying the regularity prior is the fraction of structural relaxations that completed successfully. Structures were relaxed using the LBFGS algorithm [liu1989limited] with a force tolerance of 0.05 eV/Å for Density-Functional Theory and 0.001 eV/Å for all potentials. Density-Functional Theory optimizations used a maximum of 200 steps, whilst with potentials a maximum of 20 batches of 100 steps was used with 0.001 Å of Gaussian noise added to the positions every 100 steps to aid convergence. With Density-Functional Theory and GAP the number of failures was <<1% whilst with $\sigma=0$ Å $\sim$ 11% failed. For $\sigma=0.707$ Å this dropped back below 1% and remained below 2% for all larger values.
@@ III.2 Aspirin
The effect of applying the Gaussian regularity prior was also tested on ACE potentials fit to Aspirin molecular dynamics trajectories from the rMD17 dataset [christensen2020role]. This data was chosen as it relates to an organic molecule and contains multiple chemical element, both in contrast with the single-element condensed-phased Silicon dataset. In SOAP-GAP potentials it is customary to specify a different value of $\sigma$ for each element, with larger values used for larger elements. Here, as the scaling of the basis functions $\gamma_{nl}$ depends on $\sigma/r_{\mathrm{cut}}$, see Equation 34, and we choose to scale both $\sigma$ and the cutoff $r_{\mathrm{cut}}$ in the same way for different element-pairs, $\gamma_{nl}$ has no element dependence. A convenient consequence of this is that a single value of $\sigma$, combined with the default scaling based on tabulated covalent radii, specifies the regularity prior completely. In practice, $\gamma_{nl}$ is computed using the largest cutoff present for any element-pair, so the specified value of $\sigma$ corresponds to this cutoff and the effective sigma for all other element-pairs is reduced accordingly.
The training and test sets used here contained 1000 configurations each and were obtained by drawing samples from the original trajectories at equi-spaced intervals, rather than using fully random splits. ACE potentials were then fit to this training set using a smaller basis sets, with polynomial degree=(18,18,12) for the $N=1,2$ and 3 features respectively, and a larger basis set with $N=4$ and polynomial degree=(18,18,18,12). As before, an Agnesi transformation was applied to the radial coordinate and the fits was performed using standard Tikhonov regularization with varying strengths. The best achievable force errors on the test set are tabulated in Table 1 and are shown as a function of regularization strength for the $N=4$ potentials in Figure 11. With the smaller basis, a very modest 4% reduction in the force RSME on the test set is observed when using $\sigma=0.5$ compared to no regularity prior. However, on increasing the correlation order (and number of basis functions) we see a more significant difference, with $\sigma=0.5$ reducing the test error by 27%. It is also notable that without the regularity prior, the extra basis functions cause no reduction in the test error.
Table 1: Force RMSE on the Aspirin test set for ACE potentials fit using maximum correlation order NN=3 and 4 and Gaussian regularity priors with various $\sigma$.
@@
The chronic underestimation of energy by the $\sigma=0$ potentials seen in Figure 12 suggests that using larger values of $\sigma$ may lead to more stable molecular dynamics as there appear to be fewer “holes” in the PES. To test this we took 100 configurations in the test set and attempted to perform 1 ns of MD at both 300 and 500 K. The dynamics was performed using the ase python package [larsen2017atomic] with a 0.5 fs time step and a Langevin thermostat [vanden2006second] with a friction coefficient of 3 ps-1. The simulations were terminated whenever a covalent bond present in the original molecule broke or if a new bond was formed. Thresholds for bonds breaking/forming were determined by analyzing the minimum/maximum distances present between bonded elements in the training set and then adding/subtracting an additional 0.5 Å  of tolerance. In practice the details of these criteria were found to be unimportant as any change in the number of bonds was typically accompanied by a dramatic failure of the potential where the molecule exploded. The distribution of the duration of these MD runs is shown in Figure 13 and shows that using the regularity prior led to $\sim 10\times$ longer average simulation times. To verify that this increase was not caused exclusively by over-stabilization of certain configurations the joint distribution of dihedral angles was inspected for trajectories computed with each potential. The distributions are visualized in Figure S12 and show no evidence that the trajectories with larger $\sigma$ explored less of configuration space. Furthermore, the minimum energy paths between minima were investigated using NEBs [henkelman2000climbing] and found to be in good agreement with Density-Functional Theory for all potentials, see Figure S13. As such, we conclude that applying a regularity prior provides a significant improvement to MD stability.
@@
Figure 13: Duration of MD simulations of a single Aspirin molecule before encountering a “hole” i.e. catastrophic failure. The simulations were performed at 300 and 500 K and were terminated after 1 ns if failure had not occurred.
@@
We stress that the aim of this test was not to obtain a robust potential, but instead study the effects of the regularity priors on MD stability. As such, by design we have chosen to use a relatively small amount of training data so that the MD is not stable and the effect of the regularity priors can be easily seen. To demonstrate that the MD stability can easily be improved by adding more data, we generated 10,000 configurations by applying random rotations to all 5 dihedral angles marked in Figure 12. These were then filtered down to $\sim 6500$ configurations using minimum distance constraints before farthest-point-sampling was used to select 284 new configurations to be added to the training set. The entire process was repeated and used to generate an additional 144 configurations which were added to the test set. By generating the data in this way, as opposed to sampling from the MD trajectories generated using the ACE potentials, the new configurations are completely unbiased with respect to any of the potentials, see Figure S9 for the energy and force distribution. This enlarged dataset was then refit using the same procedure as described previously. The force errors on the new test are shown in Figure S10 and, as before, applying a regularity prior leads to a significant reduction (40%) in test force RMSE. Repeating the MD simulations we see a clear improvement in MD stability, with the new durations shown in Figure 14. Finally, we also note that the additional data leads to greatly improved dihedral scans with excellent agreement compared to Density-Functional Theory seen for all potentials in Figure S11.
@@
Figure 14: Duration of MD simulations of a single Aspirin molecule before encountering a “hole” i.e. catastrophic failure, using the ACE potentials fit to the additional data generated via random dihedral rotations. The simulations were performed at 300 and 500 K and were terminated after 1 ns if failure had not occurred.

### v2

@@ III.1 Silicon
In the next test an atom in the top layer of a Si slab, a (110) surface, was moved in the $z$ direction until the atom sat just below the bottom surface of the slab due to periodic boundary conditions. The energy and z-component of the force on the target atom are shown in figure 7; as before, increasing the regularity prior leading to better agreement with the Density-Functional Theory reference, particularly for the force. Notably, using larger $\sigma$ smooths out the unphysical oscillations present from 2 Å and onwards but does not smooth out the relatively sharp feature around 0.5 Å which is present in the Density-Functional Theory reference. The same test was repeated for the (100) and (111) surfaces with the results shown in figure S3.
Refer to caption(a)(b)(c)(d)
@@
[…]
@@
Next, ACE potentials were fit to the full Silicon-GAP-18 dataset. The aforementioned tests were repeated with the new models, see figure S7, and similar improvements were seen when utilising a regularity prior, suggesting that such priors still improve performance even in a much more data-rich setting. Secondly, compression simulations were performed, wherein the low-density amorphous (LDA) structure of Si from [deringer2021origins] was gradually compressed during a Molecular Dynamics simulation by applying a pressure ramp from zero to 20 GPa. It is expected [deb2001pressure, mcmillan2005density, pandey2011pressure] that the material undergoes a sequence of phase transitions: LDA $\rightarrow$ (very high density amorphous) VHDA $\rightarrow$ polycrystalline simple-hexagonal (pc-sh). This sequence is known to be reproduced by GAP models trained on the full Silicon-GAP-18 dataset [deringer2021origins], and can serve as a challenging test for interatomic potentials [morrow2022indirect, morrow2023validate]. Indeed, it was previously shown that the nucleation of pc-sh was not consistently produced by Moment Tensor Potentials (MTPs) trained directly on the Si-GAP-18 dataset [morrow2022indirect]. Instead, consistent nucleation could only be achieved with MTP models if the training dataset was augmented with large, 1000 atom, configurations spanning the desired pressure range but labeled with a GAP model fit to the original data. This “student–teacher” model relationship was exploited to transfer the desired behaviour from the accurate, but slower, GAP model to the much faster MTP, enabling larger simulations to be performed. In the present work, we explore whether using a regularity prior might allow us to bypass the need for the addition of synthetic data and instead to achieve the desired sequence of phase transitions with ACE models fitted directly to the Silicon-GAP-18 dataset.
@@
The results of such compression simulations, carried out as in ref. [morrow2022indirect], are shown in figure 9, where the SOAP similarity to the pressure adjusted pc-sh phase is shown as a function of pressure for ACE models fit with 11 different levels of Gaussian regularity prior. With the default prior, $\sigma=0$, the dynamics become unstable during the transition from LDA $\rightarrow$ VHDA. Specifically, the model predicts spuriously large forces that accelerate atoms to velocities exceeding the integrator’s stability threshold, resulting in lost atoms and termination of the simulation. Prior to the crash, we observe the formation of chemically-unreasonable clusters of highly-coordinated Si, visible as blue patches in Figure 9B–C. These are false minima (“holes”) in the PES accessed under the influence of compression. Increasing $\sigma$ leads to a gradual increase in stability, with the $\sigma=1.0$ Å model successfully completing the full simulation. With moderate levels of regularity prior, $\sigma=1.0-2.0$ Å, nucleation of the pc-sh phase is successfully observed, indicating that a well chosen regularity prior has helped mitigate the need for the student-teacher relationship. At larger values of $\sigma$ the dynamics becomes less stable, with no successful nucleation found to occur. This worsening performance correlates with the increasing errors shown in figure S6, and implies that a well chosen $\sigma$ is one that approximately minimizes the force error on the test set.
Refer to caption(a)(b)(c)(d)(e)
@@
Figure 9: (e) Following ref. [morrow2022indirect], the average SOAP similarity to the pressure adjusted sh-Si phase is shown for 100,000 atom simulations of amorphous Si during compression from 0-20 GPa. The sharp increase around 10 GPa indicates the LDA $\rightarrow$ VHDA transition whilst the subsequent increase occurs on nucleation of the pc-sh phase. The difference in SOAP similarity compared to the GAP reference is shown below (note the break in the y-axis) whilst representative snapshots are highlighted above. (a) Atoms are colored according to local coordination number to highlight an erroneously highly coordinated cluster in the top right corner which occurs due to a hole in the potential. (b) and (d) No nucleation of the sh-Si phase is seen with some regions adopting an unusual local structure with hints of hexagonal symmetry (c) nucleation of the sh-crystal is seen with multiple differnet nucleation sites present. Similar structures are seen with $\sigma=1.41$ and 2.0 Å, with simulations remaining stable as the growing grains impinge on each other, likely inducing large local stresses.
Refer to caption(a)(b)(c)(d)(e)(f)(g)(h)(i)
@@
Figure 10: The distribution of energy, relative to the diamond structure, and volume of random structures relaxed with various energy methods are shown. The axes are linear within the central square bounded by dashed grey lines and logarithmic outside. Four illustrative structures which are all predicted to be overly stable are show in (c) consists of isolated clusters, (d) is a 3D lattice where the closest distance between any pair of atoms is 3.7Å  and (e) and (f) are both 1D “nano-wires” where some atoms have more than 4 immediate neighbours.
@@
Following this, the Silicon ACE potentials were used to perform a random structure search. The search was restricted to 8 atom unit cells with 1000 putative structures generated without symmetry constraints and a further 2000 generated using 2-4 randomly chosen symmetry operations. Minimum distance constraints of 1.7Å  were enforced between neighbouring atoms and the volume of the initial cell was drawn from a uniform distribution spanning 12-30 $\text{\AA }^{3}$/atom. Structural relaxations were carried out with 0.01 GPa of external pressure applied in lieu of a dispersion-correction scheme and a force tolerance of 0.001 eV/Å. The distribution of relaxed structures is shown in figure 10 for searches performed using Density-Functional Theory (CASTEP [clark2005first]), GAP and representative ACE models. The GAP search correctly identified the diamond structure as lowest in energy but contains a much larger proportion of low density, high energy structures than the Density-Functional Theory reference. In contrast, the ACE models all incorrectly find at least one structure to be lower in energy than diamond, with such structures all comprised of locally dense, highly coordinated, Si clusters or nano-wires. Among the ACE potentials, using a non-zero $\sigma$ significantly improves the marginal energy and volume distributions compared to the Density-Functional Theory reference and helps removes the low density, low energy structures. This result might have been expected from figure 8 where the spurious oscillations in the volume distribution are eliminated. However, we stress that a regularity prior will not remove all spurious minima in the predicted PES and indeed this is the case, as is seen for the search performed with $\sigma=2.0$ Å. Perhaps the clearest benefit of applying the regularity prior is the fraction of structural relaxations that completed successfully. Structures were relaxed using the LBFGS algorithm [liu1989limited] with a force tolerance of 0.05 eV/Å for Density-Functional Theory and 0.001 eV/Å for all potentials. Density-Functional Theory optimizations used a maximum of 200 steps, whilst with potentials a maximum of 20 batches of 100 steps was used with 0.001 Å of Gaussian noise added to the positions every 100 steps to aid convergence. With Density-Functional Theory and GAP the number of failures was <<1% whilst with $\sigma=0$ Å $\sim$ 11% failed. For $\sigma=0.707$ Å this dropped back below 1% and remained below 2% for all larger values.
@@ III.2 Aspirin
The effect of applying the Gaussian regularity prior was also tested on ACE potentials fit to Aspirin Molecular Dynamics trajectories from the rMD17 dataset [christensen2020role]. This data was chosen as it relates to an organic molecule and contains multiple chemical element, both in contrast with the single-element condensed-phased Silicon dataset. In SOAP-GAP potentials it is customary to specify a different value of $\sigma$ for each element, with larger values used for larger elements. Here, $\gamma_{nl}$ has no element dependence as the scaling of the basis functions $\gamma_{nl}$ depends on $\sigma/r_{\mathrm{cut}}$, see Equation 34, and we choose to scale both $\sigma$ and the cutoff $r_{\mathrm{cut}}$ in the same way for different element-pairs. A convenient consequence of this is that a single value of $\sigma$, combined with the default scaling based on tabulated covalent radii, specifies the regularity prior completely. In practice, $\gamma_{nl}$ is computed using the largest cutoff present for any element-pair, so the specified value of $\sigma$ corresponds to this cutoff and the effective $\sigma$ for all other element-pairs is reduced accordingly.
[…]
Table 1: Force RMSE on the Aspirin test set for ACE potentials fit using maximum correlation order NN=3 and 4 for Gaussian regularity priors with various $\sigma$.
@@
The chronic underestimation of energy by the $\sigma=0$ potentials seen in figure 12 suggests that using larger values of $\sigma$ may lead to more stable Molecular Dynamics as there appear to be fewer “holes” in the PES. To test this we took 100 configurations in the test set and attempted to perform 1 ns of MD at both 300 and 500 K. The dynamics was performed using the ase python package [larsen2017atomic] with a 0.5 fs time step and a Langevin thermostat [vanden2006second] with a friction coefficient of 3 ps-1. The simulations were terminated whenever a covalent bond present in the original molecule broke or if a new bond was formed. Thresholds for bonds breaking/forming were determined by analyzing the minimum/maximum distances present between bonded elements in the training set and then adding/subtracting an additional 0.5 Å  of tolerance. In practice the details of these criteria were found to be unimportant as any change in the number of bonds was typically accompanied by a dramatic failure of the potential where the molecule exploded. The distribution of the duration of these MD runs is shown in figure 13 and shows that using the regularity prior led to $\sim 10\times$ longer average simulation times. To verify that this increase was not caused exclusively by over-stabilization of certain configurations the joint distribution of dihedral angles was inspected for trajectories computed with each potential. The distributions are visualized in figure S12 and show no evidence that the trajectories with larger $\sigma$ explored less of the configuration space. Furthermore, the minimum energy paths between minima were investigated using NEBs [henkelman2000climbing] and found to be in good agreement with Density-Functional Theory for all potentials, see figure S11. As such, we conclude that applying a regularity prior provides a significant improvement to MD stability.
Refer to caption(a)(b)
@@
Figure 13: Duration of MD simulations of a single Aspirin molecule before encountering a “hole” i.e. catastrophic failure. The simulations were performed at (a) 300 and (b) 500 K and were terminated after 1 ns if failure had not occurred.
@@
We stress that the aim of this test was not to obtain a robust potential, but instead study the effects of the regularity priors on MD stability. As such, by design we have chosen to use a relatively small amount of training data so that the MD is not stable and the effect of the regularity priors can be easily seen. To demonstrate that the MD stability can easily be improved by adding more data, we generated 10,000 configurations by applying random rotations to all 5 dihedral angles marked in figure 12. These were then filtered down to $\sim 6500$ configurations using minimum distance constraints before farthest-point-sampling was used to select 284 new configurations to be added to the training set. The entire process was repeated and used to generate an additional 144 configurations which were added to the test set. By generating the data in this way, as opposed to sampling from the MD trajectories generated using the ACE potentials, the new configurations are completely unbiased with respect to any of the potentials, see figure S9 for the energy and force distribution. This enlarged dataset was then refit using the same procedure as described previously. The force errors on the new test set are shown in figure S10 and, as before, applying a regularity prior leads to a significant reduction (40%) in test force RMSE. Repeating the MD simulations we see a clear improvement in MD stability, with the new durations shown in figure 14. Finally, we also note that the additional data leads to greatly improved dihedral compared to Density-Functional Theory seen for all potentials in figure S13.
Refer to caption(a)(b)
@@
Figure 14: Duration of MD simulations of a single Aspirin molecule before encountering a “hole” i.e. catastrophic failure, using the ACE potentials fit to the additional data generated via random dihedral rotations. The simulations were performed at (a) 300 and (b) 500 K and were terminated after 1 ns if failure had not occurred.